OpenStudy (anonymous):

The rate at which a plant grows is constant. At the 3rd week it is 12 inches tall and at the 5th week it is 20 inches tall. If this information is plotted on a coordinate plane, what is the y-coordinate of the the point (1, y)? A) 1 B) 2 C) 4 D) 9

OpenStudy (p0sitr0n):

they ask you to find the equation of the line and then calculate the y coordinate from it. so the slope m=20-12 / 5-3=8/2=4 so y=4x+b plug in one of the points, say 3,12 12=4*3+b b=0 so y=4x, x is time in weeks, y is the height y=4*1=4 so the answer is c), 4

OpenStudy (anonymous):

Thank you very much :)